OpenStudy (stylesprettygirl):

Hey guys I really need some help on slopes can anyone help me What is the slope of the line? A. –3 B. -1/3 C.1/3 D

OpenStudy (stylesprettygirl):

If anyone ones online I could really use some help with this please

OpenStudy (stylesprettygirl):

Just a explanation would help very very much and If anyone does not mind I could use help on a few more

OpenStudy (kropot72):

To find the slope of a line through 2 points, (x0,y0) and (x1,y1), use the following formula: \[\large m=\frac{y _{1}-y _{0}}{x _{1}-x _{0}}\] Plugging in the values of the points we get: \[\large m=\frac{-3-6}{5-2}=-\frac{9}{3}\]

OpenStudy (kropot72):

The values are plugged in for you. You just need to simplify the fraction: \[\large -\frac{9}{3}\]

OpenStudy (stylesprettygirl):

Thank You so much You explained perfectly